# DGQ: Distribution-Aware Group Quantization for Text-to-Image Diffusion Models

### v1

@@ Abstract
Despite the widespread use of text-to-image diffusion models across various tasks, their computational and memory demands limit practical applications. To mitigate this issue, quantization of diffusion models has been explored. It reduces memory usage and computational costs by compressing weights and activations into lower-bit formats. However, existing methods often struggle to preserve both image quality and text-image alignment, particularly in lower-bit($<$ 8bits) quantization. In this paper, we analyze the challenges associated with quantizing text-to-image diffusion models from a distributional perspective. Our analysis reveals that activation outliers play a crucial role in determining image quality. Additionally, we identify distinctive patterns in cross-attention scores, which significantly affects text-image alignment. To address these challenges, we propose Distribution-aware Group Quantization (DGQ), a method that identifies and adaptively handles pixel-wise and channel-wise outliers to preserve image quality. Furthermore, DGQ applies prompt-specific logarithmic quantization scales to maintain text-image alignment. Our method demonstrates remarkable performance on datasets such as MS-COCO and PartiPrompts. We are the first to successfully achieve low-bit quantization of text-to-image diffusion models without requiring additional fine-tuning of weight quantization parameters.
@@ 4.1 Implementation details
Baseline and implementation details. We use two state-of-the-art methods, Q-Diffusion (Li et al., 2023a) and TFMQ-DM (Huang et al., 2024), as baselines for comparison. To ensure a fair evaluation, we employ the diffusers  [^3] https://github.com/huggingface/diffusers library for both the baselines and our method. For the text-to-image models, we conducted tests using Stable Diffusion v1.4. Unless specified otherwise, we apply 25 inference steps for computational efficiency. It should be noted that Q-Diffusion and TFMQ-DM set the attention score’s quantizer bits of Stable Diffusion to 16 bits to avoid text-image alignment degradation. However, in our implementation, to ensure a fair comparison, we set all attention score’s quantizer bits to match the activation bits.
@@ 4.2 Main results
We conducted experiments on the MS-COCO and PartiPrompts Datasets, with the results presented in Table 2. First, on the MS-COCO Dataset, all results show that our DGQ significantly outperforms previous methods. Specifically, with 8-bit activations, DGQ consistently delivers the best performance, regardless of the weight bits. The FID scores were 13.15 and 13.28 for the W8A8 and W4A8 settings, respectively, surpassing the performance of the full precision model (14.44). For text-image alignment, the CLIP score experienced only minimal drops, with decreases of 0.001 and 0.004.
For settings below 8 bits, such as 6-bit activations, previous baseline methods essentially failed to generate viable images, while DGQ successfully produced images of acceptable quality. Although there was inevitable performance degradation compared to the full precision model, our results showed significant improvement over the baseline methods. These improvements were evident in both the FID(from more than 200 to 43.66 on a 4/6 setting.) and CLIP scores(from 0.155 to 0.274 on a 4/6 setting.).
@@ 4.3 Ablation study
We performed an ablation study to analyze the impact of each component of our proposed method. The experimental results are presented in Table 3. All experiments were conducted under the W8A8 setting.

### v2

@@ Abstract
Despite the widespread use of text-to-image diffusion models across various tasks, their computational and memory demands limit practical applications. To mitigate this issue, quantization of diffusion models has been explored. It reduces memory usage and computational costs by compressing weights and activations into lower-bit formats. However, existing methods often struggle to preserve both image quality and text-image alignment, particularly in lower-bit($<$ 8bits) quantization. In this paper, we analyze the challenges associated with quantizing text-to-image diffusion models from a distributional perspective. Our analysis reveals that activation outliers play a crucial role in determining image quality. Additionally, we identify distinctive patterns in cross-attention scores, which significantly affects text-image alignment. To address these challenges, we propose Distribution-aware Group Quantization (DGQ), a method that identifies and adaptively handles pixel-wise and channel-wise outliers to preserve image quality. Furthermore, DGQ applies prompt-specific logarithmic quantization scales to maintain text-image alignment. Our method demonstrates remarkable performance on datasets such as MS-COCO and PartiPrompts. We are the first to successfully achieve low-bit quantization of text-to-image diffusion models without requiring additional fine-tuning of weight quantization parameters. Code is available at https://github.com/ugonfor/DGQ.
@@ 4.1 Implementation details
Baseline and implementation details. We use two state-of-the-art methods, Q-Diffusion (Li et al., 2023a) and TFMQ-DM (Huang et al., 2024), as baselines for comparison. To ensure a fair evaluation, we employ the diffusers  [^3] https://github.com/huggingface/diffusers library for both the baselines and our method. Unless specified otherwise, we apply 25 inference steps for computational efficiency. It should be noted that Q-Diffusion and TFMQ-DM set the attention score’s quantizer bits to 16 bits to avoid text-image alignment degradation. However, in our implementation, to ensure a fair comparison, we set all attention score’s quantizer bits to match the activation bits.
@@ 4.2 Main results
We conducted experiments on the MS-COCO and PartiPrompts datasets, with the results presented in Table 2. First, on the MS-COCO dataset, all results indicate that our DGQ significantly outperforms existing methods. Specifically, with 8-bit activations, DGQ consistently delivers the best performance, regardless of the weight bit width. The FID scores were 13.15 and 13.28 for the W8A8 and W4A8 settings, respectively, surpassing the performance of the full-precision model (14.44). For text-image alignment, the CLIP score experienced only minimal drops (less than 0.005).
For the 6-bit activation setting, baseline methods failed to generate images, whereas DGQ successfully produced images of acceptable quality. Although there was inevitable performance degradation compared to the full-precision model, our results showed significant improvement over the baseline methods. These improvements were evident in both the FID score (reducing from over 200 to 43.66 in the W4A6 setting) and the CLIP score (increasing from 0.155 to 0.274 in the W4A6 setting).
@@ 4.3 Ablation study
We performed an ablation study to analyze the impact of each component of our proposed method. The experimental results are presented in Table 3.
@@ 5 Conclusion
Acknowledgements
We thank all CVML members for valuable feedbacks. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075 Artificial Intelligence Graduate School Program(KAIST), No.2021-0-02068 Artificial Intelligence Innovation Hub, No. RS-2024-00457882, National AI Research Lab Project), IITP grant funded by the Korea government(MSIT) and KEIT grant funded by the Korea government(MOTIE) (No. 2022-0-00680, 2022-0-01045), and Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the MSIP (NRF-2022R1A2C3011154, RS-2023-00219019).